# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v3

@@ 1 Introduction
Experimental results show that $\text{D}^{3}\text{RE}$ improves robustness and efficiency in downstream tasks such as density ratio estimation, mutual information estimation, and density estimation. Figure 1 illustrates a comparison of interpolation strategies among DI, DBI, DDBI, and DSBI, with light blue points representing intermediate samples drawn from $q_{0}$ and $q_{1}$. Our proposed methods (DBI, DDBI, and DSBI) enable broader exploration of alternative trajectories, producing intermediate distributions with larger support compared to DI, consistent with Theorem 3.2 and Corollary 3.3 below.
@@
Figure 1: Trajectory sets comparison among DI, DBI, DDBI and DSBI. Our methods yield broader trajectory sets, with intermediate distributions exhibiting wider support than those of DI. The entropically regularized transport losses for subfigures (a-d) are $44.17$, $31.14$, $31.15$, and $8.26$, respectively (See Equation 17 for details.). Lower loss indicates increased path diversity.
@@
  • We incorporate guidance mechanisms to improve interpolant quality: GD improves the stability of time score functions at boundary (Theorem 3.4,Corollary 3.8), while OTR leads to more accurate (Theorem 3.7) and efficient (Figure 8) estimation.
@@ 2.2 Denoising Diffusion Bridge Model
Denoising diffusion bridge models (Zhou et al., 2024a) overcome this limitation by simulating stochastic processes that interpolate between paired distributions with $\mathbf{X}_{1}$ as endpoints. These processes are derived from the SDE in Eq. (6) via Doob’s $h$-transform (Doob & Doob, 1984)
$\mathrm{d}\mathbf{X}_{t}=\tilde{\mathbf{f}}(\mathbf{X}_{t},t,\mathbf{X}_{1})\mathrm{d}t+g(t)\mathrm{d}\mathbf{W}_{t},$ (7)
where $\tilde{\mathbf{f}}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\mathbf{f}(\mathbf{X}_{t},t)+g^{2}(t)\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})$ denotes the shifted drift and $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{1}\mid\mathbf{X}_{t})$ is the $h$ function representing the gradient of the log transition kernel from time $t$ to 1. The process explicitly depends on $\mathbf{X}_{1}$ and, given $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, it forms a diffusion bridge with transition kernel $q_{t}(\mathbf{x}\mid\mathbf{X}_{0},\mathbf{X}_{1})$. A special case of the diffusion bridge, termed the Brownian bridge, arises under the conditions $\mathbf{f}(\mathbf{X}_{t},t):=0,g(t):=1$ and $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1}):=\frac{\mathbf{X}_{1}-\mathbf{X}_{t}}{1-t}$. The Brownian bridge is defined as the solution to $\mathrm{d}\mathbf{B}_{t}=\frac{\mathbf{B}_{1}-\mathbf{B}_{t}}{1-t}\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}$ and its transition kernel is given by $q(\mathbf{B}_{t}\mid\mathbf{B}_{0},\mathbf{B}_{1})=\mathcal{N}((1-t)\mathbf{B}_{0}+t\mathbf{B}_{1},t(t-1)\mathbf{E}_{d})$.
@@ 3.1 Support-chasm Problem
Deterministic Interpolant. We summarize the interpolants used in prior works, such as TRE and DRE-$\infty$, to deterministic interpolant (DI), defined as
@@
Prior methods, such as TRE (Rhodes et al., 2020) and DRE-$\infty$ (Choi et al., 2022), are specific cases of DI distinguished by their choices of $\alpha_{t}$ and $\beta_{t}$ (see Section B.1).
@@
Diffusion Bridge Interpolant. To mitigate the support-chasm, we introduce a Brownian bridge $\{\mathbf{B}_{t}\}_{t\in[0,1]}$, leading to the diffusion bridge interpolant (DBI)
@@
Since $\mathbf{B}_{t}$ is a Gaussian process with zero mean and variance $t(1-t)$, the transition kernel of the DBI is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(\mathbf{I}(\mathbf{x}_{0},\mathbf{x}_{1},t),t(1-t)\gamma^{2}\mathbf{E}_{d})$, leading to an equivalent form of DBI
@@
See Section A.1 for detailed derivation. This result shows that DBI covers a larger or equal region of the space compared to DI, providing theoretical justification for its ability to mitigate the support-chasm problem in $\text{D}^{3}\text{RE}$.
@@
See Section A.2 for details. This generalizes the support expansion result to entire paths, implying that DBI, by injecting noise, explores a broader set of trajectories than DI. As a result, it provides better coverage of the interpolation space, enhancing robustness across diverse distributions.
@@ 3.2 Gaussian Dequantization for Mollified Time Score
Let $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ be a DI defined in Equation 8. Under A.1 and A.2, the time score for any $t\in(0,1)$ satisfies
@@
where $L$ is the Lipschitz constant in A.2. Moreover, if $L<1$, this lower bound diverges to infinity
@@
Proofs can be found in Section A.3.
Dequantified Diffusion Bridge Interpolant. To stabilize the time score near the boundary, we introduce Gaussian dequantization (GD) by adding controlled perturbations to boundary samples, effectively handling the boundary densities and resulting uniformly bounded time score across $[0,1]$ (see Corollary 3.8 for details).
@@
Here, the additional term $(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon$ reflects the effect of GD, yielding smoother interpolation. As a result, the transition kernel of the DDBI, $q_{t}^{\prime}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})$, is given by $\mathcal{N}(\mathbf{I}(\mathbf{x}_{0},\mathbf{x}_{1},t),\left(t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon\right)\mathbf{E}_{d})$, which shares the same mean trajectory as DBI (see Equation 10).
@@
See Section A.4 for detailed derivation. Proposition 3.5 confirms that the dequantified density ratio $r^{\prime}(\mathbf{x})$ is a uniform approximation of $r(\mathbf{x})$ for sufficiently small $\varepsilon$.
@@ 3.3 Optimal Transport Rearrangement
where $\Pi(q_{0}^{\prime},q_{1}^{\prime})$ is the set of all probability paths with marginals $q_{0}^{\prime}$ and $q_{1}^{\prime}$, and $\mathcal{H}(\pi)$ is the entropy of $\pi$. The coefficient $2\gamma^{2}$ is regularization factor (details in Section B.2).
We apply the scalable Sinkhorn algorithm (Cuturi, 2013) to mini-batches $\{\mathbf{x}_{0}^{n}\}_{n=1}^{N}{\sim}q_{0}^{\prime}(\mathbf{x})$ and $\{\mathbf{x}_{1}^{n}\}_{n=1}^{N}{\sim}q_{1}^{\prime}(\mathbf{x})$, obtaining rearranged pairs $\{\hat{\mathbf{x}}_{0}^{n},\hat{\mathbf{x}}_{1}^{n}\}_{n=1}^{N}{\sim}\pi_{2\gamma^{2}}$. This procedure is termed as optimal transport rearrangement (OTR) for convenience.
Dequantified Schrödinger Bridge Interpolant. Applying OTR followed by DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ yields the dequantified Schrödinger bridge interpolant (DSBI)
$\mathbf{X}_{t}=\mathbf{I}(\hat{\mathbf{X}}_{0}^{\prime},\hat{\mathbf{X}}_{1}^{\prime},t)+\sqrt{t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon}\mathbf{Z}_{t},$ (18)
where $\mathbf{I}(\hat{\mathbf{X}}_{0}^{\prime},\hat{\mathbf{X}}_{1}^{\prime},t)=(1-t)\hat{\mathbf{X}}_{0}^{\prime}+t\hat{\mathbf{X}}_{1}^{\prime}$.
@@
See Section A.5 for proof. This result suggests that DSBI, as a principled integration of DDBI and OTR, implicitly solves the SB problem and provides a minimum-cost stochastic interpolation between $q_{0}^{\prime}$ and $q_{1}^{\prime}$.
@@
Consider the DDBI and DSBI with coefficients $\alpha_{t}=1-t$, $\beta_{t}=t$. Let $\pi\in\Pi(q_{0}^{\prime},q_{1}^{\prime})$ be an arbitrary coupling for DDBI and $\pi_{2\gamma^{2}}$ the entropically OT coupling for DSBI. Define the error functional $\mathcal{E}:=\big{|}\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\,\mathrm{d}t-\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\big{|}$. Then, the upper bound of $\mathbb{E}_{\pi_{2\gamma^{2}}}[\mathcal{E}_{\text{DSBI}}^{2}]$ is less than or equal to that of $\mathbb{E}_{\pi}[\mathcal{E}_{\text{DDBI}}^{2}]$.
@@
See Section A.6 and Section A.7 for detailed derivations of Theorem 3.7 and Corollary 3.8. These results establish that both DDBI and DSBI admit a bounded time-integrated time score, $\mathbb{E}_{q_{t}^{\prime}}[|\partial_{t}\log q_{t}^{\prime}|]$, in contrast to the divergent lower bound exhibited by DI (see Theorem 3.4). Moreover, it provides a formal justification for the error reduction achieved by DSBI through the coupling $\pi_{2\gamma^{2}}$.
@@ 3.4 Dequantified Diffusion Bridge DRE
For the DDBI, $r^{\prime}(\mathbf{x})$ can be approximated effectively using a neural network, as formulated in Definition 3.9. See Section A.8 for detailed derivation.
@@
where $\partial_{t}\left[\lambda(t)s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)\right]=\lambda(t)\partial_{t}s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)+\lambda^{\prime}(t)s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)$, $\partial_{t}s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)$ and $\lambda^{\prime}$ denote the time derivative of the time score model and weighting function, respectively. The first two terms enforce the boundary conditions. $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ differ only by a constant $C$ independent of $\boldsymbol{\theta}$. In practice, for stable and effective training, the joint score-matching loss is implemented, as described in Section C.2.
Estimating Target Log Density Ratio. Given the optimal parameters $\boldsymbol{\theta}^{\star}$ obtained by minimizing $\mathcal{L}3$, the log density ratio at any point $\mathbf{x}$ can be estimated as
@@
[…]
@@
[…]
  $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi_{2\gamma^{2}}$
[…]

### v5

@@ 1 Introduction
Experimental results show that $\text{D}^{3}\text{RE}$ improves robustness and efficiency in downstream tasks such as density ratio estimation, mutual information estimation, and density estimation. Fig. 1 illustrates a comparison of interpolation strategies among DI, DBI, DDBI, and DSBI, with light blue points representing intermediate samples drawn from $q_{0}$ and $q_{1}$. Our proposed methods (DBI, DDBI, and DSBI) enable broader exploration of alternative trajectories, producing intermediate distributions with larger support compared to DI, consistent with Theorem 3.2 and Corollary 3.3 below.
@@
Figure 1: Trajectory sets comparison among DI, DBI, DDBI and DSBI. Our methods yield broader trajectory sets, with intermediate distributions exhibiting wider support than those of DI. The entropically regularized transport losses for subfigures (a-d) are $44.17$, $31.14$, $31.15$, and $8.26$, respectively (See Eq. 17 for details.). Lower loss indicates increased path diversity.
@@
  • We incorporate guidance mechanisms to improve interpolant quality: GD improves the stability of time score functions at boundary (Theorem 3.4,Corollary 3.8), while OTR leads to more stable (Theorem 3.7) and efficient (Fig. 8) estimation.
@@ 2.2 Denoising Diffusion Bridge Model
Denoising diffusion bridge models (Zhou et al., 2024a) overcome this limitation by simulating stochastic processes that interpolate between paired distributions with $\mathbf{X}_{1}$ as endpoints. These processes are derived from the SDE in Eq. (6) via Doob’s $h$-transform (Doob & Doob, 1984),
$\mathrm{d}\mathbf{X}_{t}=[\mathbf{f}(\mathbf{X}_{t},t)+g^{2}(t)\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})]\mathrm{d}t+g(t)\mathrm{d}\mathbf{W}_{t},$ (7)
where $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{1}\mid\mathbf{X}_{t})$ is the $h$ function representing the gradient of the log transition kernel from time $t$ to 1. The process explicitly depends on $\mathbf{X}_{1}$ and, given $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, it forms a diffusion bridge with transition kernel $q_{t}(\mathbf{x}\mid\mathbf{X}_{0},\mathbf{X}_{1})$. A special case of the diffusion bridge, termed the Brownian bridge, arises under the conditions $\mathbf{f}(\mathbf{X}_{t},t):=0,g(t):=1$ and $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1}):=\frac{\mathbf{X}_{1}-\mathbf{X}_{t}}{1-t}$. The Brownian bridge is defined as the solution to $\mathrm{d}\mathbf{B}_{t}=\frac{\mathbf{B}_{1}-\mathbf{B}_{t}}{1-t}\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}$ and its transition kernel is given by $q_{t}(\mathbf{b}\mid\mathbf{B}_{0},\mathbf{B}_{1})=\mathcal{N}((1-t)\mathbf{B}_{0}+t\mathbf{B}_{1},t(1-t)\mathbf{E}_{d})$.
@@ 3.1 Support-chasm Problem
Deterministic Interpolant.
We summarize the interpolants used in prior works, such as TRE and DRE-$\infty$, to deterministic interpolant (DI), defined as
@@
Prior methods, such as TRE (Rhodes et al., 2020) and DRE-$\infty$ (Choi et al., 2022), are specific cases of DI distinguished by their choices of $\alpha_{t}$ and $\beta_{t}$ (see Sec. B.1).
@@
Diffusion Bridge Interpolant.
To mitigate the support-chasm, we introduce a Brownian bridge $\{\mathbf{B}_{t}\}_{t\in[0,1]}$, leading to the diffusion bridge interpolant (DBI)
@@
Since $\mathbf{B}_{t}$ is a Gaussian process with zero mean and variance t(1t)t(1-t), i.e., $\mathbf{B}_{t}\sim\mathcal{N}(\mathbf{0},t(1-t)\mathbf{E}_{d})$, the transition kernel of the DBI can be derived follows from Eq. 9 as $q_{t}(\mathbf{x}\mid\mathbf{X}_{0},\mathbf{X}_{1})=\mathcal{N}(\mathbf{I}(\mathbf{X}_{0},\mathbf{X}_{1},t),t(1-t)\gamma^{2}\mathbf{E}_{d})$. By applying the reparameterization trick, DBI admits the equivalent form
@@
See Sec. A.1 for detailed derivation. This result shows that DBI covers a larger or equal region of the space compared to DI, providing theoretical justification for its ability to mitigate the support-chasm problem in $\text{D}^{3}\text{RE}$.
@@
See Sec. A.2 for details. This generalizes the support expansion result to entire paths, implying that DBI, by injecting noise, explores a broader set of trajectories than DI. As a result, it provides better coverage of the interpolation space, enhancing robustness across diverse distributions.
@@ 3.2 Gaussian Dequantization for Mollified Time Score
Let $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ be a DI defined in Eq. 8. Under Assumption A.1 and Assumption A.2, the time score for any $t\in(0,1)$ satisfies
@@
where $L$ is the Lipschitz constant in Assumption A.2. Moreover, if $L<1$, this lower bound diverges to infinity
@@
Proofs can be found in Sec. A.3.
Dequantified Diffusion Bridge Interpolant.
To stabilize the time score near the boundary, we introduce Gaussian dequantization (GD) by adding controlled perturbations to boundary samples, effectively handling the boundary densities and resulting uniformly bounded time score across $[0,1]$ (see Corollary 3.8 for details).
@@
Here, the additional term $(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon$ reflects the effect of GD, yielding smoother interpolation. As a result, the transition kernel of the DDBI, $q_{t}^{\prime}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})$, is given by $\mathcal{N}(\mathbf{I}(\mathbf{x}_{0},\mathbf{x}_{1},t),\left(t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon\right)\mathbf{E}_{d})$, which shares the same mean trajectory as DBI (see Eq. 10).
@@
See Sec. A.4 for detailed derivation. Proposition 3.5 confirms that the dequantified density ratio $r^{\prime}(\mathbf{x})$ is a uniform approximation of $r(\mathbf{x})$ for sufficiently small $\varepsilon$.
@@ 3.3 Optimal Transport Rearrangement
where $\Pi(q_{0}^{\prime},q_{1}^{\prime})$ is the set of all probability paths with marginals $q_{0}^{\prime}$ and $q_{1}^{\prime}$, and $\mathcal{H}(\pi)$ is the entropy of $\pi$. The coefficient $2\gamma^{2}$ is regularization factor (details in Sec. B.2).
We apply the scalable Sinkhorn algorithm (Cuturi, 2013) to mini-batches $\{\left(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime}\right)_{n}\}_{n=1}^{N}\sim q_{0}^{\prime}\times q_{1}^{\prime}$, obtaining rearranged sample pairs $\{\left(\hat{\mathbf{x}}_{0}^{\prime},\hat{\mathbf{x}}_{1}^{\prime}\right)_{n}\}_{n=1}^{N}{\sim}\pi_{2\gamma^{2}}$. For convenience, we refer to this procedure as optimal transport rearrangement (OTR), which preserves the marginals and only changes the sample pairing.
Dequantified Schrödinger Bridge Interpolant.
Applying OTR followed by DDBI yields the dequantified Schrödinger bridge interpolant (DSBI)
$\hat{\mathbf{X}}_{t}^{\prime}=\mathbf{I}(\hat{\mathbf{X}}_{0}^{\prime},\hat{\mathbf{X}}_{1}^{\prime},t)+\sqrt{t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon}\mathbf{Z}_{t},$ (18)
where $\alpha_{t}=1-t$ and $\beta_{t}=t$ are fixed.
@@
See Sec. A.5 for proof. This result suggests that DSBI, as a principled integration of DDBI and OTR, implicitly solves the SB problem and provides a minimum-cost stochastic interpolation between $q_{0}^{\prime}$ and $q_{1}^{\prime}$.
@@
Consider the DDBI and DSBI with $\alpha_{t}=1-t$, $\beta_{t}=t$. Let $\pi\in\Pi(q_{0}^{\prime},q_{1}^{\prime})$ be any coupling for DDBI, and $\pi_{2\gamma^{2}}$ the entropically regularized OT coupling for DSBI. Then, for all $t\in[0,1]$, the variance of the time score under DSBI is no greater than that under DSBI, i.e.,
$\mathrm{Var}^{\text{DSBI}}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\leq\mathrm{Var}^{\text{DDBI}}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime}).$ (20)
@@
See Sec. A.6 and Sec. A.7 for detailed derivations of Theorem 3.7 and Corollary 3.8. These results establish that both DDBI and DSBI admit a bounded time-integrated time score, $\mathbb{E}_{q_{t}^{\prime}}[|\partial_{t}\log q_{t}^{\prime}|]$, in contrast to the divergent lower bound exhibited by DI (see Theorem 3.4). Moreover, it provides a formal justification for the error reduction achieved by DSBI through the coupling $\pi_{2\gamma^{2}}$.
@@ 3.4 Dequantified Diffusion Bridge DRE
For the DDBI, $r^{\prime}(\mathbf{x})$ can be approximated effectively using a neural network, as formulated in Definition 3.9. See Sec. A.8 for detailed derivation.
@@
where $\partial_{t}\left[\lambda(t)s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)\right]=\lambda(t)\partial_{t}s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)+\lambda^{\prime}(t)s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)$, $\partial_{t}s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)$ and $\lambda^{\prime}$ denote the time derivative of the time score model and weighting function, respectively. The first two terms enforce the boundary conditions. $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ differ only by a constant $C$ independent of $\boldsymbol{\theta}$. In practice, for stable and effective training, the joint score-matching loss is implemented, as described in Sec. C.2.
Estimating Target Log Density Ratio.
Given the optimal parameters $\boldsymbol{\theta}^{\star}$ obtained by minimizing $\mathcal{L}3$, the log density ratio at any point $\mathbf{x}$ can be estimated as
@@
$\mathbf{x}^{\prime}_{0}\leftarrow\mathbf{x}_{0}+\mathbf{z}_{\varepsilon},\mathbf{x}^{\prime}_{1}\leftarrow\mathbf{x}_{1}+\mathbf{z}_{\varepsilon}$ $\%\ \text{GD}$~~~~
$\mathbf{x}_{t}^{\prime}\leftarrow\alpha_{t}\mathbf{x}_{0}^{\prime}+\beta_{t}\mathbf{x}_{1}^{\prime}+\sqrt{t(1-t)\gamma^{2}}\mathbf{z}$ $\%\ \text{DBI}\,$~~
@@
$\mathbf{x}^{\prime}_{0}\leftarrow\mathbf{x}_{0}+\mathbf{z}_{\varepsilon},\mathbf{x}^{\prime}_{1}\leftarrow\mathbf{x}_{1}+\mathbf{z}_{\varepsilon}$ $\%\ \text{GD}$~~~~
$\pi_{2\gamma^{2}}\leftarrow\mathsf{Sinkhorn}(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime},2\gamma^{2})$ $\%\ \text{OTR}$~~
$(\hat{\mathbf{x}}_{0}^{\prime},\hat{\mathbf{x}}_{1}^{\prime})\sim\pi_{2\gamma^{2}}$
$\hat{\mathbf{x}}^{\prime}\leftarrow\alpha_{t}\hat{\mathbf{x}}_{0}^{\prime}+\beta_{t}\hat{\mathbf{x}}_{1}^{\prime}+\sqrt{t(1-t)\gamma^{2}}\mathbf{z}$ $\%\ \text{DBI}\,$~~